OpenStudy (anonymous):

Frank is on his way home in his car. He has driven 15 miles so far, which is three-fifths of the way home. What is the total length of his drive?

OpenStudy (texaschic101):

3/5 of x = 15 3/5x = 15 x = 15/(3/5) x = 15 * 5/3 x = 75/3 x = 25 check... 3/5 * 25 = 75/5 = 15 miles 2/5 * 25 = 50/5 = 10 miles -------------------------add 5/5 = 25 miles The total length of the drive is 25 miles.